# Schwinger dark matter production

### v1

@@ II.1 Generating dark electric fields during inflation
In order to form a classical dark electric field which can source Schwinger production during and after inflation, the vector bosons must be produced in a configuration with a very large occupation number and a coherence wavelength atleast of order the horizon, $H^{-1}$. Here we describe how to generate such a dark electric field during inflation. Mechanisms for producing a vector field during inflation can be constructed in a number of ways. One possibility is to consider a minimal coupling to the metric via a vector mass term Graham et al. (2016); Ahmed et al. (2020). Alternatively, one can consider non-minimal couplings between the inflaton and the gauge field. The two most common such couplings are ones between (a function of the) the inflaton and the kinetic term of the gauge field $I(\phi)FF$ Shakeri et al. (2019); Nakayama (2020); Nakai et al. (2020) or the dual field strength $I(\phi)F\widetilde{F}$ Bastero-Gil et al. (2019, 2022); Bastero-Gil and Manso (2022).
@@
Note the conductivity enters as a correction to the Hubble damping term in Eq. (18). As stated in Eq. (6), we work in the regime $\bar{\sigma}\ll 1$, so we can neglect the effect of the conductivity on the transverse vector mode equation of motion. Taking $\xi$ to be time-independent, the solution can be found in terms of Whittaker functions using the WKB approximation as shown in Bastero-Gil et al. (2022) for instance, to which we refer the reader for details. From Eq. (18) we see that when $k/a\lesssim 2\xi H$, one of the transverse modes becomes tachyonic (which we take to be $A_{+}$) and exponentially enhanced (large occupation number) which we can associate with particle production of the dark vector Bastero-Gil et al. (2019, 2022). In our regime of interest, where we can neglect backreaction effects while still generating enough particle production, $\xi$ typically takes values between 1 and 9 at the end of inflation. Thus exponential particle production occurs when the wavelength of the dark vector is comparable to the size of the horizon. This generates a classical background dark electric field with a coherence length of order the inverse Hubble scale. This dark electric field can then act as a source for Schwinger pair production during and after inflation.
@@
To estimate how large the dark electric field is we can compute the energy density contained in the electric field. Since the parameter $\xi$ increases with time, as $\dot{\phi}$ increases and $H$ decreases, it is largest at the end of inflation and in principle depends on time. In Bastero-Gil et al. (2022) the effect of this time dependence was taken into account in numerical solutions of Eq. (18) and shown that, while the time dependance can effect the shape and peak of the energy density power spectrum, the average energy density is sufficiently well approximated using analytic solutions to obtain order of magnitude estimates of the dark matter parameter space. Using the WKB approximation to solve Eq. (18) one finds Bastero-Gil et al. (2019, 2022),
$\rho_{E}$$\equiv$ $\langle\rho_{E}\rangle=\frac{1}{2}\langle|E_{+}|^{2}\rangle=\frac{1}{2}\langle\frac{1}{a^{2}}|\partial_{t}A_{+}|^{2}\rangle$ (21)
$\approx$ $10^{-4}\frac{e^{2\pi\xi}}{\xi^{3}}H^{4}.$
@@
Here $\rho_{E}\equiv\langle\rho_{E}\rangle$ represents the spacial average of the total dark electric field energy density, where most of the energy is stored as the one in the dark magnetic field is suppressed by $\rho_{B}=\rho_{E}/\xi^{2}$. As shown in Bastero-Gil et al. (2022), this result is in decent agreement with the numerical result which takes into account the time variation of $\xi$. Taking $\xi$ and $H$ constant, we see from Eq. (21) that the magnitude of the physical dark electric field sourced by the rolling inflaton remains constant. This is a reasonable approximation when we are close to the end of inflation as $\xi$ and Hubble change little in the last few e-folds. Since the energy density is exponentially sensitive to $\xi$ which increases towards the end of inflation, as the slow roll parameter increases, the vast majority of the dark electric field is produced at the end of inflation. Hence, in the last few e-folds of inflation we obtain a roughly constant dark electric field of order the Hubble horizon and we can apply Eq. (8) and Eq. (9) to compute the energy density of the Schwinger produced dark charged particles.
@@
To produce particles during inflation it is in general necessary to break conformal invariance. We emphasize that here we consider a single source of conformal breaking generated by the axion like coupling between the inflaton and the dark photon, $\alpha/f$ in Eq. (17) which introduces a new scale. As discussed above, we take the dark charged particles $\chi$ to be conformally coupled by neglecting their mass during inflation, $m_{\chi}\ll H$ (and by taking $\xi_{R}=1/6$ in Eq. (2) for the scalar case). This ensures there is no purely gravitational particle production Ford (1987); Kuzmin and Tkachev (1999) and we can isolate the the particle production due to the inflationary Schwinger effect.
II.2 Conductivity and weak electric force regime
@@
So far we have used the conductivity $\sigma$ to parametrize and quantify the energy density of the particles produced via the Schwinger effect. Such an effect in flat spacetime depends on the “electric force” ($g_{D}E$) as well as on the mass of the charged particle and leads to an exponential suppression $\propto{\rm exp}[-m_{\chi}^{2}/|g_{D}E|]$. In de-Sitter space expansion introduces one more parameter in the Hubble scale $H$ and non exponentially suppressed regimes become possible. In this section we give a brief account of how the conductivity is expected to behave in the regime we are interested in. For this we rely on previous studies of the Schwinger effect in a de-Sitter background Kobayashi and Afshordi (2014); Hayashinaka et al. (2016); Banyeres et al. (2018).
@@
The strong electric force regime where $g_{D}E>H^{2}$ was studied in Bastero-Gil et al. (2023) which showed that this condition also implies $m_{\chi}>H$. In this regime dark Schwinger production dominates over the exponentially suppressed (with $m_{\chi}/H$) contribution arising from purely gravitational effects. Furthermore, the Schwinger effect exhibits the same behaviour in de-Sitter space as in flat space Kobayashi and Afshordi (2014); Banyeres et al. (2018). However, in the weak force regime, no Schwinger production occurs in flat space while in de-Sitter space it can still occur and exhibits various non-intuitive features Hayashinaka et al. (2016); Banyeres et al. (2018). Here we examine whether dark Schwinger production can also generate enough dark matter in the weak electric force regime where,
$|g_{D}E|\ll H^{2}\,.$ (22)
As discussed above, in order to ensure it is not dominated by (purely) gravitational effects Ford (1987); Kuzmin and Tkachev (1999), we must also be in the conformal limit where $m_{\chi}/H_{\rm end}\ll 1$ during inflation. (and $\xi_{R}=1/6$ for the scalar case).
As our dark electric field has a magnitude typically much larger than $H^{2}$, given the exponential enhancement in Eq. (21), we must have a weak gauge coupling,
$g_{D}\ll\frac{H^{2}}{E}=10^{2}\frac{\xi^{3/2}}{e^{\pi\xi}}.$ (23)
To see how the conductivity is related to the gauge coupling let us consider the case of scalar dark charged particles Kobayashi and Afshordi (2014); Banyeres et al. (2018). In the weak electric force regime the conductivity is given by Kobayashi and Afshordi (2014),
$\sigma=g_{D}^{2}\frac{H}{24\pi^{2}}$ $\left[\ln\left(\frac{\xi_{R}R}{H^{2}}\right)+\frac{16\pi}{3}\frac{\mu_{0}(-1+\mu_{0}^{2})}{\sin(2\pi\mu_{0})}\right.$
$\left.-\psi\left(\frac{1}{2}+\mu_{0}\right)-\psi\left(\frac{1}{2}-\mu_{0}\right)\right]\,,$ (24)
where we have defined,
$\mu_{0}^{2}=\frac{9}{4}-\frac{\xi_{R}R}{H^{2}}\,,$ (25)
with $\psi$ the digamma function and we have taken $m_{\chi}=0$ in Eq. (2). Setting $\xi_{R}=1/6$ for the conformally coupled scalar and recalling that $R=12H^{2}$ in de-Sitter gives,
$\bar{\sigma}=\frac{\sigma}{H}=\frac{g_{D}^{2}}{24\pi^{2}}\left[\ln 2+\frac{2}{3}+2\gamma_{E}\right]\approx 10^{-2}g_{D}^{2}\,,$ (26)
where $\gamma_{E}$ is the Euler gamma constant. Note we find a similar order of magnitude estimate for the fermion case as well Hayashinaka et al. (2016). We see that the conductivity is $\propto g_{D}^{2}$ so a small gauge coupling $g_{D}\ll 1$ guarantees a small $\sigma$ in this case. In the next section we will see that the cosmic evolution imposes bounds on $g_{D}$ stronger than Eq. (23).
@@ III Cosmic evolution and dark
$\rho^{\rm{end}}_{E}\equiv\rho_{A}(T_{\rm RH})\approx 10^{-4}\frac{e^{2\pi\xi_{\rm end}}}{\xi_{\rm end}^{3}}\,H_{\rm end}^{4}\,.$ (28)
@@
Here $g_{*}(T_{\rm RH})$ denotes the number of relativistic degrees of freedom which we fix to $g_{*}(T_{\rm RH})\sim 100$, restricting ourselves to reheating temperatures above the electroweak scale. We see Eq. (30) also defines the relation between the value of Hubble at reheating $H_{\rm RH}$, which we introduced in Eq. (11), and $H_{\rm end}$ at the end of inflation,
$H_{\rm RH}=\epsilon_{R}^{2}H_{\rm end}\,.$ (31)
At reheating the dark vector and $\chi$ have energy densities,
$\rho_{A}(T_{\rm RH})$ $\equiv\rho_{E}^{\rm RH}=\rho_{E}^{\rm end}\,,$ (32)
$\rho_{\chi}^{\rm RH}$ $=\rho_{\chi}^{\rm end}\,,$ (33)
@@
$\rho_{\chi}^{\rm RH}\ll\rho_{A}(T_{\rm RH})\ll\rho_{R}(T_{\rm RH})\ll\rho^{\rm{end}}_{I}\,.$ (34)
We assume that the dark and visible sectors are decoupled and that the dark sector does not thermalize with itself. While at reheating $\rho_{\chi}$ starts as the smallest energy density, it grows during the radiation domination era due to continued Schwinger production (see Eq. (II)). Furthermore, when the momentum redshifts below $m_{\chi}$, at which point Schwinger production becomes highly suppressed, the energy density in $\chi$ will begin to redshift like matter as $a^{-3}$ and eventually becomes dominant after matter radiation equality. For the mechanism considered here, the dark vector can be massless, in which case its energy density continues redshifting like radiation and remains negligibly subdominant throughout its cosmic evolution. In what follows however, we consider a massive vector, in which case it eventually becomes non-relativistic and contributes to the dark matter abundance also. The massless case is recovered by setting $m_{A}=0$.
@@
$q_{A}(T_{\rm RH})\equiv\frac{k}{a_{\rm RH}}=\frac{k}{a_{\rm end}}\simeq H_{\rm end}\gg m_{A}\,.$ (35)
@@
$\rho_{A}(T_{0})=\left(\frac{T_{0}}{T_{RH}}\right)^{3}\left(\frac{m_{A}}{H_{\rm end}}\right)\rho_{A}(T_{\rm RH})\,.$ (40)
As shown in Bastero-Gil et al. (2019, 2022), the dark vector can account for the entirety of the dark matter over a very large range of dark vector masses $\sim\mu eV-10\,TeV$. Furthermore, due to large density fluctuations, today the vector dark matter has a clumpy like nature on physical scales $\sim cm-100\,km$ Bastero-Gil et al. (2022).
Next we turn to the evolution of $\rho_{\chi}$. After reheating the energy density of $\chi$ evolves as described by Eq. (II), with $n=4$ (relativistic) and $\delta=2$ (radiation),
$\rho_{\chi}(T)=\frac{2\bar{\sigma}_{\rm RH}\rho_{E}^{\rm RH}}{2-\alpha}\left[\left(\frac{T}{T_{\rm RH}}\right)^{2+\alpha}\,-\,\frac{(2+\alpha)}{4}\left(\frac{T}{T_{\rm RH}}\right)^{4}\right].~{}~{}$ (41)
Here $\alpha$ parametrizes the possible redshift scaling of the conductivity $\sigma=\sigma_{\rm RH}(T/T_{\rm RH})^{\alpha}$ as defined in Eq. (11)). If $20\leq\alpha<$ 2 the first term in Eq. (41) dominates as $T$ decreases and $\rho_{\chi}(T)$ redshifts slower than radiation. The case $\alpha=2$ corresponds to the conductivity redshifting like Hubble and results in,
$\rho_{\chi}(T)=\frac{\bar{\sigma}_{\rm RH}\rho_{E}^{\rm RH}}{2}\left(\frac{T}{T_{\rm RH}}\right)^{4}\left[1-\ln\left(\frac{T}{T_{\rm RH}}\right)^{4}\right]\,.$ (42)
Thus it redshifts almost like radiation apart from a logarithmic correction. If $\alpha>2$ the last term in Eq. (41) dominates and the redshift is like radiation. We see that only if the conductivity redshifts slower than Hubble do we have a significant contribution to $\rho_{\chi}$ during radiation domination. Thus, if $\alpha\geq 2$ we can neglect Schwinger production after reheating and just follow the evolution of $\rho_{\chi}$ that was produced during inflation.
The next step is to determine when $\chi$ becomes non-relativistic which requires tracking its momentum. During inflation $\chi$ is produced at the (negligible) expense of the energy density of the dark electric field. As $\rho_{E}$ peaks at a wavelength the size of the horizon, it is reasonable to expect that $\chi$ inherits a power spectrum peaked at the same wavelength. Thus, the typical momentum of $\chi$ at reheating is the same as that of the vector,
@@
This evolves in time according to,
$\frac{dq_{\chi}}{dt}+Hq_{\chi}=g_{D}E(t)\,.$ (44)
The momentum redshifts due to the expansion of the universe, but at the same time it tends to increase due to the dark electric force accelerating $\chi$. The physical dark electric field $E(t)$ decreases in time. Again we switch to temperature to track the evolution which gives,
$E(T)=\sqrt{2\rho_{E}(T)}=\sqrt{2\rho_{E}^{\rm RH}}\left(\frac{T}{T_{\rm RH}}\right)^{2}\,,$ (45)
and Eq. (44) becomes,
$\frac{dq_{\chi}}{dT}-\frac{q_{\chi}}{T}=-\frac{g_{D}}{T}\frac{\sqrt{2\rho_{E}^{\rm RH}}}{H_{\rm RH}}\,,$ (46)
which has a solution of the form,
$q_{\chi}(T)=q_{\chi}(T_{\rm RH})\frac{T}{T_{\rm RH}}+g_{D}\frac{\sqrt{2\rho_{E}^{\rm RH}}}{H_{\rm RH}}\left(1-\frac{T}{T_{\rm RH}}\right)\,.~{}~{}~{}$ (47)
The momentum must decrease with decreasing temperature if $\chi$ is to become non-relativistic. This implies a maximum dark gauge coupling $g_{D}<H_{\rm end}H_{\rm RH}/\sqrt{2\rho_{E}^{\rm RH}}$ and reflects the fact that the gauge coupling must be weak enough to ensure that momentum redshifting dominates over the effect of the acceleration due to the dark electric field. However, there are more stringent bounds on the gauge coupling coming from requiring $\chi$ to become non-relativistic by the time of matter radiation equality (MRE) at temperature $T_{\rm MRE}\sim{\rm eV}$. We can solve for the temperature at which $\chi$ becomes non-relativistic using $q_{\chi}(\overline{T}_{\chi})=m_{\chi}$ and Eq. (47) which gives,
$\overline{T}_{\chi}=\frac{H_{\rm RH}m_{\chi}-g_{D}\sqrt{2\rho_{E}^{\rm RH}}}{H_{\rm end}H_{\rm RH}-g_{D}\sqrt{2\rho_{E}^{\rm RH}}}T_{\rm RH}\,.$ (48)
Thus, we have the condition $T_{\rm MRE}<\overline{T}_{\chi}<T_{\rm RH}$ which translates into the bound on the gauge coupling
$g_{D}$ $<$ $\frac{\epsilon_{R}^{2}H_{\rm end}^{2}}{E}\left(\frac{m_{\chi}}{H_{\rm end}}-\frac{T_{\rm MRE}}{T_{\rm RH}}\right)$ (49)
$=$ $50\sqrt{2}\frac{\xi_{\rm end}^{3/2}}{e^{\pi\xi_{\rm end}}}\epsilon_{R}^{2}\left(\frac{m_{\chi}}{H_{\rm end}}-\frac{T_{\rm MRE}}{T_{\rm RH}}\right)\,,$
together with the condition,
$\frac{T_{\rm MRE}}{T_{\rm RH}}<\frac{m_{\chi}}{H_{\rm end}}\ll 1\,,$ (50)
to keep $g_{D}$ positive which will be the case over all of the parameter space we consider. In fact Eq. (49) also holds when $m_{\chi}>H_{\rm end}$, but since we are working in the conformal limit we limit ourselves to $m_{\chi}<H_{\rm end}$. Note that Eq. (49) gives a much stronger upper bound on $g_{D}$ compared to Eq. (23), due to the further suppression proportional to $\epsilon_{R}^{2}(m_{\chi}/H_{\rm end})$. The bound in Eq. (49) can push the gauge coupling to very small values in some regions of parameter space and we restrict ourselves to the region in which we satisfy the weak gravity conjecture Arkani-Hamed et al. (2007),

### v2

@@ II.1 Generating dark electric fields during inflation
In order to form a classical dark electric field which can source Schwinger production during and after inflation, the vector bosons must be produced in a configuration with a very large occupation number and a coherence wavelength at least of order the horizon, $H^{-1}$. Here we describe how to generate such a dark electric field during inflation. Mechanisms for producing a vector field during inflation can be constructed in a number of ways. One possibility is to consider a minimal coupling to the metric via a vector mass term Graham et al. (2016); Ahmed et al. (2020). Alternatively, one can consider non-minimal couplings between the inflaton and the gauge field. The two most common such couplings are ones between (a function of) the inflaton and the kinetic term of the gauge field $I(\phi)FF$ Shakeri et al. (2019); Nakayama (2020); Nakai et al. (2020) or the dual field strength $I(\phi)F\widetilde{F}$ Bastero-Gil et al. (2019, 2022); Bastero-Gil and Manso (2022).
@@
Note the conductivity enters as a correction to the Hubble damping term in Eq. (12). As stated in Eq. (6), we work in the regime $\bar{\sigma}\ll 1$, so we can neglect the effect of the conductivity on the transverse vector mode equation of motion. Taking $\xi$ to be time-independent, the solution can be found in terms of Whittaker functions using the WKB approximation as shown in Bastero-Gil et al. (2022) for instance, to which we refer the reader for details. From Eq. (12) we see that when $k/a\lesssim 2\xi H$, one of the transverse modes becomes tachyonic (which we take to be $A_{+}$) and exponentially enhanced which we can associate with particle production of the dark vector Bastero-Gil et al. (2019, 2022).
In our regime of interest, where we can neglect backreaction effects while still generating enough particle production, $\xi$ takes values between 1 and 9 at the end of inflation. Thus exponential particle production occurs when the wavelength of the dark vector is comparable to the size of the horizon. This generates a classical background dark electric field with a coherence length of order the inverse Hubble scale. This dark electric field can then source Schwinger pair production during inflation.
@@
To estimate how large the dark electric field is we can compute the energy density contained in the electric field. Since the parameter $\xi$ increases with time, as $\dot{\phi}$ increases and $H$ decreases, it is largest at the end of inflation and in principle depends on time. In Bastero-Gil et al. (2022) the effect of this time dependence was taken into account in numerical solutions of Eq. (12) and shown that, while the time dependance can affect the shape and peak of the energy density power spectrum, the average energy density is sufficiently well approximated using analytic solutions to obtain order of magnitude estimates of the dark matter parameter space. Using the WKB approximation to analytically solve Eq. (12) one finds Bastero-Gil et al. (2019, 2022),
$\rho_{E}$$\equiv$ $\langle\rho_{E}\rangle=\frac{1}{2}\langle\frac{1}{a^{2}}|\partial_{t}A_{+}|^{2}\rangle$ (15)
$=$ $\frac{1}{2}\langle|E_{+}|^{2}\rangle\approx 10^{-4}\frac{e^{2\pi\xi}}{\xi^{3}}H^{4}.$
Here $\rho_{E}\equiv\langle\rho_{E}\rangle$ represents the spacial average of the total dark electric field energy density. An accompanying dark magnetic field is also produced, but it is suppressed by $\rho_{B}=\rho_{E}/\xi^{2}$. This magnetic field is automatically produced parallel to the electric field (see Eq. (46) in the Appendix) and would thus lead to a small enhancement of the Scwhinger effect Fujita et al. (2022), but for the rough estimates of the parameter space we make here it can be neglected.
@@
As shown in Bastero-Gil et al. (2022), the analytic result in Eq. (15) is in decent agreement with the numerical result which takes into account the time variation of $\xi$ with time. Taking $\xi$ and $H$ constant, we see from Eq. (15) that the magnitude of the physical dark electric field sourced by the rolling inflaton remains constant. This is a reasonable approximation when we are close to the end of inflation as $\xi$ and Hubble change little in the last few e-folds. Since the energy density is exponentially sensitive to $\xi$ which increases towards the end of inflation, as the slow roll parameter increases, the vast majority of the dark electric field is produced at the end of inflation. Hence, in the last few e-folds of inflation we obtain a roughly constant dark electric field of order the Hubble horizon and we can apply Eq. (8) and Eq. (9) to compute the energy density of the Schwinger produced dark charged particles.
@@
To produce particles during inflation it is in general necessary to break conformal invariance. We emphasize that here we consider a single source of conformal breaking generated by the axion like coupling between the inflaton and the dark photon, $\alpha/f$ in Eq. (11) which introduces a new scale. As discussed above, we take the dark charged particles $\chi$ to be conformally coupled by neglecting their mass during inflation, $m_{\chi}\ll H$ (and by taking $\xi_{R}=1/6$ in Eq. (2) for the scalar case). This ensures there is no purely gravitational particle production Ford (1987); Kuzmin and Tkachev (1999) and we can isolate the particle production due to the inflationary Schwinger effect.
II.2 Conductivity &\&& weak/strong electric force
@@
So far we have used the conductivity $\sigma$ to parametrize and quantify the energy density of the particles produced via the Schwinger effect. Such an effect in flat spacetime depends on the “electric force” ($g_{D}E$) as well as on the mass of the charged particle and leads to an exponential suppression $\propto{\rm exp}[-m_{\chi}^{2}/|g_{D}E|]$. In de-Sitter space expansion introduces one more parameter in the Hubble scale $H$ and non exponentially suppressed regimes become possible. In this section we give a brief account of how the conductivity is expected to behave in the $|g_{D}E|\ll H^{2}$ (weak) and $g_{D}E\gg H^{2}$ (strong) electric force regimes. For this we rely on previous studies of the Schwinger effect in a de-Sitter background Kobayashi and Afshordi (2014); Hayashinaka et al. (2016); Banyeres et al. (2018).
The conductivity can be computed in terms of the two dimensionless ratios Kobayashi and Afshordi (2014); Hayashinaka et al. (2016); Banyeres et al. (2018),
$\lambda=\frac{g_{D}E}{H^{2}},\quad\overline{m}_{\chi}=\frac{m_{\chi}}{H},$ (16)
with which one can study different limits. The strong electric force regime is defined via the limit $\lambda\gg 1$. In this regime the Schwinger effect exhibits the same behaviour in de-Sitter space as in flat space and leads to the dimensionless conductivity Kobayashi and Afshordi (2014); Banyeres et al. (2018),
$\bar{\sigma}\equiv\frac{\sigma}{H}=\frac{g_{D}^{2}}{12\pi^{3}}\lambda e^{-\pi\overline{m}_{\chi}^{2}/\lambda}~{}~{}(\lambda\gg 1).$ (17)
This applies to scalars, while for fermions the conductivity is larger by a factor of 2 Hayashinaka et al. (2016). When $\lambda>\overline{m}_{\chi}$ we see the conductivity is not exponentially suppressed even if $m_{\chi}>H$. This implies that particle production can occur during inflation via the Schwinger effect even if purely gravitational effects, which are exponentially suppressed with $m_{\chi}/H$, are negligible. This case was studied in Bastero-Gil et al. (2023) which examined inflationary Schwinger production of superheavy dark matter with mass $m_{\chi}>H_{\rm end}$ to which we refer the interested reader.
Here we are interested in light dark matter where $\overline{m}_{\chi}<1<\lambda$. In the $\overline{m}_{\chi}\gg\lambda$ regime $\overline{\sigma}$ simplifies to,
$\bar{\sigma}=\frac{g_{D}^{2}}{12\pi^{3}}\lambda~{}~{}(\lambda\gg 1\gg\overline{m}_{\chi}).$ (18)
This implies that even in the massless (conformal) limit, where again purely gravitational effects are neglibile Ford (1987); Kuzmin and Tkachev (1999) we can have particle production during inflation due to the Schwinger effect. As we’ll see below, this opens the possibility of generating light scalar and/or fermion dark matter during inflation which is typically difficult with only purely gravitational production. Since we will only consider $\overline{m}_{\chi}<1$ in this study, Eq. (18) is a sufficient approximation for present purposes.
@@
In the weak force regime where $\lambda<1$, no Schwinger production occurs in flat space while in de-Sitter space it can occur and exhibits various non-intuitive features Hayashinaka et al. (2016); Banyeres et al. (2018). Here we examine whether dark Schwinger production can also generate enough dark matter in the weak electric force regime. In the weak electric force regime the conductivity is given by Kobayashi and Afshordi (2014); Banyeres et al. (2018),
$\bar{\sigma}=\frac{\sigma}{H}$ $=$ $\frac{g_{D}^{2}}{24\pi^{2}}\left[\ln 2+\frac{2}{3}+2\gamma_{E}\right]$ (19)
$\approx$ $10^{-2}g_{D}^{2}\,~{}~{}(\lambda\ll 1),$
where, in order to ensure (purely) gravitational effects are negligible Ford (1987); Kuzmin and Tkachev (1999), we have set $m_{\chi}=0$ and $\xi_{R}=1/6$ in Eq. (2) for a conformally coupled scalar and $R=12H^{2}$ in de-Sitter. The last equality holds approximately for both dark charged fermions and scalars Kobayashi and Afshordi (2014); Hayashinaka et al. (2016); Banyeres et al. (2018) and will be useful when exploring the dark matter relic density parameter space as a function of $g_{D}$.
@@ III Cosmic evolution and dark
$\rho^{\rm{end}}_{E}=\frac{1}{2}E^{2}\approx 10^{-4}\frac{e^{2\pi\xi_{\rm end}}}{\xi_{\rm end}^{3}}\,H_{\rm end}^{4}\,.$ (21)
@@
Here $g_{*}(T_{\rm RH})$ denotes the number of relativistic degrees of freedom which we fix to $g_{*}(T_{\rm RH})\sim 100$, restricting ourselves to reheating temperatures above the electroweak scale. We see Eq. (23) also defines the relation between the value of Hubble at reheating $H_{\rm RH}$ and $H_{\rm end}$ at the end of inflation,
$H_{\rm RH}=\epsilon_{R}^{2}H_{\rm end}\,,$ (24)
as well as the reheating temperature $T_{\rm RH}$. At reheating the dark vector and $\chi$ have energy densities,
$\rho_{A}(T_{\rm RH})$ $=$ $\rho_{E}^{\rm end}\,,$ (25)
$\rho_{\chi}(T_{\rm RH})$ $=$ $\rho_{\chi}^{\rm end}\,,$ (26)
@@
$\rho^{\rm{end}}_{\chi}\ll\rho^{\rm{end}}_{E}\ll\rho_{R}(T_{\rm RH})\ll\rho^{\rm{end}}_{I}\,.$ (27)
We assume that the dark and visible sectors are decoupled and that the dark sector does not thermalize with itself. While at reheating $\rho_{\chi}$ starts as the smallest energy density, once the momentum of $\chi$ redshifts below its mass $\rho_{\chi}$ will begin to redshift like matter as $a^{-3}$ and eventually becomes dominant after matter radiation equality.
For the mechanism considered here, the dark vector could in principle be massless, in which case its energy density continues redshifting like radiation and remains negligibly subdominant throughout its cosmic evolution. However, if the dark vector is massless limits from structure formation and plasma instabilities Lasenby (2020), which are stronger than those coming from triaxiality of dark matter hallows Lasenby (2020); Agrawal et al. (2017), put severe constraints on the dark gauge couplings and would rule out much of the dark matter parameter space for inflationary Schwinger production. These constraints can be avoided if the dark vector has a mass larger than $m_{A}\gtrsim 10^{-10}\,eV$ so at some point during its cosmic evolution, it must obtain a mass. Thus in what follows, we consider a massive dark vector, in which case it eventually becomes non-relativistic and contributes to the dark matter relic abundance although for the masses we consider this contribution will be negligible. The massless case is recovered by setting $m_{A}=0$.
@@
$q_{A}(T_{\rm RH})\equiv\frac{k}{a_{\rm end}}\simeq H_{\rm end}\gg m_{A}\,.$ (28)
@@
$\rho_{A}(T_{0})=\left(\frac{T_{0}}{T_{RH}}\right)^{3}\left(\frac{m_{A}}{H_{\rm end}}\right)\rho^{\rm{end}}_{E}\,.$ (33)
where we have used Eq. (25). As shown in Bastero-Gil et al. (2019, 2022), the dark vector can account for the entirety of the dark matter over a very large range of dark vector masses $\sim\mu$eV $-10$ TeV. Furthermore, due to large density fluctuations, today the vector dark matter has a clumpy like nature on physical scales $\sim$ cm $-100$ km Bastero-Gil et al. (2022). In what follows we will always consider $m_{A}\ll m_{\chi}<H_{\rm end}$ so that the dark photon contributes negligibly to the dark matter relic abundance which is dominated by $\chi$.
Turning now to the evolution of $\rho_{\chi}$ we must first determine the momentum of $\chi$ at the end of inflation (and instantaneous reheating) as done for the dark vector in Eq. (35). For the dark electrons which are Schwinger produced during inflation there are competing effects between damping due to expansion and the acceleration due to the background dark electric field. We can write the evolution of the physical momentum of $\chi$ as,
$q(\tau)_{\chi}=q(\tau_{0})_{\chi}\,e^{-N}+\lambda H(1-e^{-N}),$ (34)
where $q(\tau_{0})$ is the momentum at production in conformal time which depends on the electric field production mechanism and $N$ counts the number e-folds from $\tau_{0}$. The first term is the usual damping term due to expansion. The second term is due to the presence of the background electric field which accelerates the electrons. The net effect is that during inflation $\chi$ very quickly reaches a terminal momentum of $\lambda H$. This holds generically for any mechanism which produces an electric field during inflation that sources Schwinger production.
However, for the dark electric field production mechanism considered above and examined in Bastero-Gil et al. (2022), the dark electric field is dominantly produced just at the end of inflation. In this case the dark electrons have not had much time to either redshift or get accelerated by the electric field. Thus taking $N\approx 0$ and assuming the dark electrons are produced with momentum comparable to the dark photons, at the end of inflation (and instantaneous reheating) we again have for the initial momentum,
@@
A precise determination of $q_{\chi}(T_{\rm RH})$ requires numerical study, but could potentially further expand the dark matter parameter space examined below.
As we detail in the Appendix, once inflation ends the electric field quickly begins to oscillate Bastero-Gil et al. (2022) and no longer accelerates the dark electrons efficiently. Thus as the universe expands and the temperature decreases the momentum for the dark electrons redshifts as in Eq. (29) for the dark photon, but with initial momentum given in Eq. (35). Thus for the energy density of $\chi$ today we can follow the same steps as for the evolution of the dark photon and immediately write,
$\rho_{\chi}(T_{0})=\left(\frac{T_{0}}{T_{RH}}\right)^{3}\left(\frac{m_{\chi}}{H_{\rm end}}\right)\frac{\overline{\sigma}}{2}\rho^{\rm{end}}_{E}\,,$ (36)
where we have used Eq. (22) and Eq. (25) for $\rho^{\rm{end}}_{\chi}$. This analysis is sensible only if the dark sector does not thermalize with the visible sector. Thermalization with the visible sector is easily avoided by taking the dark sector to be secluded. This implies that a possible kinetic mixing between visible and dark $U(1)$ has to be small.
III.1 Constraints on dark gauge coupling
The above analysis assumes the different dark sector components redshift independently, thus we also need to ensure that the dark sector does not thermalize within itself. This requires that the rate of interactions between the dark charged particles and the dark photon is always slower than Hubble which we can estimate as follows. While $\chi$ is relativistic its number density is given by $n_{\chi}(T)\simeq\rho_{\chi}(T)/q_{\chi}(T)$ and the 2-to-2 cross section, mediated by a dark photon, can be estimated as $\sigma_{2\to 2}\sim g_{D}^{4}/q^{2}_{\chi}(T)$. The scattering rate is $\Gamma=n_{\chi}(T)\sigma_{2\to 2}$. Requiring that $\Gamma(T)<H(T)$ until $\chi$ becomes non-relativistic at a temperature $\overline{T}_{\chi}=\frac{m_{\chi}}{H_{\rm end}}T_{\rm RH}$ leads to the condition $g_{D}^{4}\overline{\sigma}<\epsilon_{R}^{2}\overline{m}_{\chi}(H_{\rm end}^{2}/E)^{2}$ where $\overline{\sigma}$ is the conductivity given in Eq. (17) and Eq. (19) for the $\lambda>1$ and $\lambda<1$ regimes respectively.
In the strong electric force limit $\lambda>1$ the no thermalization in the dark sector condition leads to,
$g_{D}\lesssim 2\left(\frac{\overline{m}_{\chi}\epsilon_{R}^{2}}{(E/H_{\rm end}^{2})^{3}}\right)^{1/7}~{}~{}(\lambda>1).$ (37)
An additional upper bound on $g_{D}$ comes from requiring no backreaction from Schwinger production which could destroy the electric field and which implies $\overline{\sigma}<1$, but this leads to a weaker bound than Eq. (37). A lower bound on $g_{D}$ comes from imposing $\lambda>1$,
$g_{D}>\frac{1}{E/H_{\rm end}^{2}}~{}~{}(\lambda>1).$ (38)
In the weak electric force limit the strongest upper bound comes from imposing the condition $\lambda<1$,
$g_{D}<\frac{1}{E/H_{\rm end}^{2}}.$ (39)
An additional lower bound in both cases arises if we demand we satisfy the weak gravity conjecture Arkani-Hamed et al. (2007),